# Artificial Intelligence in Creative Industries: Advances Prior to 2025

### v3

@@ 3.7.2 Performance and main challenges
Although significant progress has been made in the past few years in visual quality assessment, including new models and training methodology, challenges remain, including limited model generalization and high computational complexity.
@@
4 Closing Thoughts and Future of AI in Creativity
This paper has presented a comprehensive review of current AI technologies and their applications that have emerged in recent years. Generative methods have driven a rapid growth in AI usage, particularly in the creative sector, significantly advancing the state of the art across various creative applications such as content creation, information extraction and analysis, content enhancement and data compression.
Through these applications, generative AI has not only broadened creative possibilities but has also reduced the manual effort and time traditionally associated with the production pipeline, allowing for greater creative experimentation and quicker production cycles. As this technology advances, it promises to unlock even more sophisticated capabilities in the creative industry. However, creative technologists, artists and other users must adapt, learn to use, and build these tools effectively and safely.
@@ 4.1 Challenges for AI in the Creative Sector
One of the primary challenges for artists engaging with modern generative AI and LLMs is the lack of consistent, controllable output. These models operate via stochastic sampling from high-dimensional latent spaces, meaning that identical prompts can yield different results across runs. This unpredictability makes it difficult for artists to achieve and iterate toward a precise creative vision. Although prompt engineering has emerged as a technique to guide model behavior, it requires technical knowledge and iterative refinement, which may not align with the intuitive or exploratory approaches common in artistic practice.
@@
Moreover, there is a fundamental tension between the structured nature of current AI pipelines and the nonlinear, often improvisational workflows of creative disciplines. Many generative tools were originally designed for tasks like software development, content automation, or optimization [419], and are ill-suited for open-ended, exploratory creation. Artists typically work in cycles of ideation, experimentation, and revision—processes that demand fluid, real-time interaction and control, which existing AI systems struggle to support. These limitations point to a gap in current AI design: a need for systems that not only generate high-quality content but also adapt to the iterative, interpretive nature of artistic production. One possible approach addressed to these challenges is a shift toward top-down creative workflows, where artists define high-level concepts, themes, or goals via text prompts before refining specific outputs. This approach helps align AI-generated results with artistic intent, offering a degree of control over inherently stochastic systems.
@@
Speaking at the World Government Summit in Dubai in 2024, [^56] https://blogs.nvidia.com/blog/world-governments-summit/ NVIDIA CEO Jensen Huang argued that with rapid advancements in AI, learning to code may become less essential for newcomers to the tech sector. He envisioned a future where traditional programming could be replaced by more intuitive AI-driven tools, thereby automating complex tasks and enhancing productivity—particularly for artists without coding expertise. While this perspective remains debated, it highlights the potential for AI to become more accessible within creative fields. However, AI-assisted coding tools are insufficient for creative practitioners, as artistic workflows tend to be unstructured and rely on domain-specific data. Artists must hence turn to techniques such as fine-tuning pre-trained models, few-shot learning, or domain adaptation—methods that are powerful yet typically inaccessible without machine learning expertise.
@@
There is also broader concerns persist regarding the long-term impact of AI on the creative industries, particularly with the potential emergence of artificial general intelligence (AGI). Envisioned by organizations like OpenAI, DeepMind, and Anthropic, AGI could surpass human cognitive abilities, raising ethical and existential questions about the role of human agency in artistic expression.
@@ 4.2 Ethical Issues, Fakes and Bias
Advancements in generative AI, exemplified by models like Sora and Gemini 1.5 Pro, provoke ethical concerns and societal implications. These models, capable of generating highly realistic content, escalate the risk of misuse, through malicious deepfakes and misinformation. We are now in a situation where AI results transcend the uncanny valley, further complicating matters and challenging perceptions of authenticity. For example, the artist Miles Astray demonstrated that even authentic photographs could be mistaken for AI-generated images. His real photograph ‘F L A M I N G O N E’ won both the jury’s award and the people’s choice award in the AI category of the 1839 Awards. His aim wasto highlight the ethical dilemmas inherent in AI, suggesting that the benefits of discussing AI’s ethical implications could surpass the ethical concerns related to viewer deception [^57] https://www.milesastray.com/news/newsflash-reclaiming-the-brain.
While democratizing AI tools no-doubt presents opportunities to transform creative processes and workflows, it also necessitates robust regulatory frameworks to safeguard privacy and ownership. For example, deepfake technologies stimulate significant concerns about the spread of misinformation and other malicious uses. Efforts to detect and identify increasingly realistic deepfakes are thus as important as the generative methods used to produce them. These must however be accompanied by increased media literacy, and policies that address the ethical and legal implications.
Diversity and representation is a key issue when using AI tools. Unified Concept Editing [54] has been proposed as a basis for image generation in digital mediums. This aims to ensure the production of safe content with diverse representation, reducing gender and racial biases. Hallucination in generative AI (the production of outputs that are not faithful representations of reality but instead contain imagined or unrealistic elements) are a further cause of concern. These undermine trust in AI processes and can be due to limitations in the training data, biases in the model architecture or imperfections in the optimization process. Hallucinations associated with LLMs are one of the issues highlighted by the UK Government [1], alongside bias, regurgitation of private data, difficulties with multi-step tasks and challenges in interpreting black-box processes.
@@ 4.3 The future of AI technologies
A substantial amount of data is essential for training AI models in order to achieve high performance and good generalisation. Major companies such as Google, Meta, and NVIDIA, with their respective models: BERT, Segment Anything, and Canvas, dominate this space, benefiting from leveraged resources to gather data and process it to train sophisticated models. However, in November 2024, Bloomberg reported that OpenAI, Anthropic, and Google are all experiencing relatively slow growth in the performance of their AI models, with one of the key challenges being training data [^60] https://www.bloomberg.com/news/articles/2024-11-13/openai-google-and-anthropic-are-struggling-to-build-more-advanced-ai.
@@
LLMs excel in applications involving complex tasks, advanced reasoning, data analysis, and understanding context. However, these models typically require high computational resources or cloud computing for development, operation and fine-tuning. A new trend emerging in alongside LLMs is the development of Small language models (SLMs), such as Phi-3 by Microsoft [^61] https://azure.microsoft.com/en-us/blog/introducing-phi-3-redefining-whats-possible-with-slms/. SLMs offer promising solutions for regulated industries and sectors encountering scenarios where high-quality results are essential while keeping data ’on-site’. Their potential is particularly relevant when deploying more capable SLMs on smartphones and other mobile devices, allowing them to operate ‘at the edge’ without relying on cloud connectivity. Recent highly successful platforms, such as DeepSeek-V3 [420] and Qwen2.5-Max [421], are based on Mixture-of-Experts (MoE) models, which tackle complex problems by dividing them into simpler sub-tasks, each handled by a specialized “expert.”
@@
Despite the evident advancements in AI, current models still struggle with tasks requiring planning or deep reasoning and are prone to errors when encountering unexpected data. This, in turn, reduces the confidence of users and trust in the results. AI algorithms can learn through reinforcement learning, but this process often identifies the best outcome as an anomaly rather than the norm. Yann LeCun, Professor at NYU and Chief AI Scientist at Meta, noted that while LLMs show a degree of comprehension in processing and generating text, their understanding lacks depth, often leading to results that defy common sense [^62] https://twitter.com/ylecun/status/1728496457601183865. He advocates for self-supervised learning as a pivotal future direction for AI, emphasizing its potential to derive insights from unlabeled data. Concurrently, Andrew Ng, Adjunct Professor at Stanford University and Founder of DeepLearning.AI, sees iterative AI agentic workflows [^63] https://www.youtube.com/watch?v=sal78ACtGTc as a key advancement for enhancing AI tool capabilities through an interactive approach by AI agents. These workflows involve autonomous agents that interactively learn from experience, understand natural language, and execute tasks on behalf of users.

### v4

@@ 3.7.2 Performance and main challenges
Although significant progress has been made in the past few years in visual quality assessment, including new models and training methodologies, challenges remain, including limited model generalization and high computational complexity.
@@
4 Closing Thoughts: The Future of AI in Creative Applications
This paper has presented a comprehensive review of current AI technologies and their creative industries applications that have emerged in recent years. Generative methods have driven a rapid growth in AI usage, particularly in the creative sector, significantly advancing the state of the art across various applications such as content creation, information extraction and analysis, content enhancement and data compression.
Through these applications, generative AI has not only broadened creative possibilities, but has also reduced the manual effort and time traditionally associated with the production pipeline, allowing for greater creative experimentation and more rapid and agile production cycles. As this technology advances, it promises to unlock even more sophisticated capabilities. However, creative technologists, artists and other users must adapt, learn to use, and build these tools effectively and safely.
@@ 4.1 Challenges for AI in the Creative Sector
One of the primary challenges for artists engaging with modern generative AI and LLMs is the lack of consistent, controllable outputs. These models operate via stochastic sampling from high-dimensional latent spaces, meaning that identical prompts can yield different results across runs. This unpredictability can make it challenging for artists to achieve, or iterate toward, a precise creative vision. Although prompt engineering has emerged as a technique to guide model behavior, it requires technical knowledge and iterative refinement, which may not align with the intuitive or exploratory approaches common in artistic practice.
@@
Moreover, there can be a fundamental tension between the structured nature of current AI pipelines and current production, often improvisational, workflows used in creative disciplines. Many generative tools were originally designed for tasks like software development, content automation, or optimization [419], and are less well suited for open-ended, exploratory creation. Artists typically work in cycles of ideation, experimentation, and revision—processes that demand fluid, real-time interaction and control, which existing AI systems struggle to support. These limitations point to a gap in current AI design: a need for systems that not only generate high-quality content but also adapt to the iterative, interpretive nature of artistic production. One possible approach to address these challenges is a reinforcement of top-down creative workflows, where artists define high-level concepts, themes, or goals via text prompts before refining specific outputs. This approach helps align AI-generated results with artistic intent, offering a degree of control over inherently stochastic systems.
@@
Speaking at the World Government Summit in Dubai in 2024, [^56] https://blogs.nvidia.com/blog/world-governments-summit/ NVIDIA CEO Jensen Huang argued that, with rapid advancements in AI, learning to code may become less essential for newcomers to the tech sector. He envisioned a future where traditional programming could be replaced by more intuitive AI-driven tools, thereby automating complex tasks and enhancing productivity—particularly for artists without coding expertise. While this perspective remains debated, it highlights the potential for AI to become more accessible within creative fields, not just in coding but across areas such as VFX and virtual production. But to achieve this, AI-assisted coding tools must be better integrated into creative workflows. Creators must exploit techniques such as fine-tuning pre-trained models, few-shot learning, or domain adaptation—methods that are powerful yet typically inaccessible without machine learning expertise.
@@
There are also broader concerns that persist regarding the long-term impact of AI on the creative industries, particularly with the potential emergence of artificial general intelligence (AGI). Envisioned by organizations like OpenAI, DeepMind, and Anthropic, AGI could surpass human cognitive abilities, raising ethical and existential questions about the role of human agency in artistic expression.
@@ 4.2 Ethical Issues, Fakes and Bias
Advancements in generative AI, exemplified by models like Sora and Gemini 1.5 Pro, provoke ethical concerns and have societal implications. WHile their applications, with appropriate permission, can be beneficial and entertaining, these models, because they are capable of generating highly realistic content, escalate the risk of misuse through malicious deepfakes and misinformation. We are now in a situation where AI results transcend the uncanny valley, further complicating matters and challenging perceptions of authenticity. For example, the artist Miles Astray demonstrated that even authentic photographs could be mistaken for AI-generated images. His real photograph ‘F L A M I N G O N E’ won both the jury’s award and the people’s choice award in the AI category of the 1839 Awards. His aim was to highlight the ethical dilemmas inherent in AI, suggesting that the benefits of discussing AI’s ethical implications could surpass the ethical concerns related to viewer deception [^57] https://www.milesastray.com/news/newsflash-reclaiming-the-brain.
While democratizing AI tools no doubt presents opportunities to transform creative processes and workflows, it also necessitates robust regulatory frameworks to safeguard privacy and ownership. For example, deepfake technologies stimulate significant concerns about the spread of misinformation and other malicious uses. Efforts to detect and identify increasingly realistic deepfakes are thus as important as the generative methods used to produce them. These must however be accompanied by increased media literacy, and policies that address the ethical and legal implications.
Diversity and representation is a key issue when using AI tools. Unified Concept Editing [54] has been proposed as a basis for image generation in digital mediums. This aims to ensure the production of safe content with diverse representation, reducing gender and racial biases. Hallucination in generative AI (the production of outputs that are not faithful representations of reality but instead contain imagined or unrealistic elements) is a further cause of concern. These undermine trust in AI processes and can be due to limitations in the training data, biases in the model architecture or imperfections in the optimization process. Hallucinations associated with LLMs are one of the issues highlighted by the UK Government [1], alongside bias, regurgitation of private data, difficulties with multi-step tasks and challenges in interpreting black-box processes.
@@ 4.3 The future of AI technologies
A substantial amount of data is essential for training AI models in order to achieve high performance and good generalization. Major companies such as Google, Meta, and NVIDIA, with their respective models: BERT, Segment Anything, and Canvas, dominate this space, benefiting from leveraged resources to gather data and process it to train sophisticated models. However, in November 2024, Bloomberg reported that OpenAI, Anthropic, and Google are all experiencing relatively slow growth in the performance of their AI models, with one of the key challenges being training data [^60] https://www.bloomberg.com/news/articles/2024-11-13/openai-google-and-anthropic-are-struggling-to-build-more-advanced-ai.
@@
LLMs excel in applications involving complex tasks, advanced reasoning, data analysis, and understanding context. However, these models typically require high computational resources or cloud computing for development, operation and fine-tuning. A new trend emerging alongside LLMs is the development of Small language models (SLMs), such as Phi-3 by Microsoft [^61] https://azure.microsoft.com/en-us/blog/introducing-phi-3-redefining-whats-possible-with-slms/. SLMs offer promising solutions for regulated industries and sectors encountering scenarios where high-quality results are essential while keeping data ’on-site’. Their potential is particularly relevant when deploying more capable SLMs on smartphones and other mobile devices, allowing them to operate ‘at the edge’ without relying on cloud connectivity. Recent highly successful platforms, such as DeepSeek-V3 [420] and Qwen2.5-Max [421], are based on Mixture-of-Experts (MoE) models, which tackle complex problems by dividing them into simpler sub-tasks, each handled by a specialized “expert.”
@@
Despite evident advancements in AI, current models still struggle with tasks requiring planning or deep reasoning and are prone to errors when encountering unexpected data. This, in turn, reduces the confidence of users and trust in the results. AI algorithms can learn through reinforcement learning, but this process often identifies the best outcome as an anomaly rather than the norm. Yann LeCun, Professor at NYU and Chief AI Scientist at Meta, noted that while LLMs show a degree of comprehension in processing and generating text, their understanding lacks depth, often leading to results that defy common sense [^62] https://twitter.com/ylecun/status/1728496457601183865. He advocates for self-supervised learning as a pivotal future direction for AI, emphasizing its potential to derive insights from unlabeled data. Concurrently, Andrew Ng, Adjunct Professor at Stanford University and Founder of DeepLearning.AI, sees iterative AI agentic workflows [^63] https://www.youtube.com/watch?v=sal78ACtGTc as a key advancement for enhancing AI tool capabilities through an interactive approach by AI agents. These workflows involve autonomous agents that interactively learn from experience, understand natural language, and execute tasks on behalf of users.